Air and fuel enter a furnace used for home heating. The air has an enthalpy of 320 kJ/kg and the fuel has an enthalpy of 43027 kJ/kg. The gases leaving the furnace has an enthalpy of 616 kJ/kg. There are 17 kgair/kgfuel, water circulated through the wall furnace receiving heat. The hose required 17.02 kW of heat, what is the fuel consumption per day. (41 kg/day)
1
Expert's answer
2020-03-16T12:17:41-0400

Here in question it is given the enthalpy of air as ha=320 KJ/kg and the enthalpy of fuel is

hf=43027 kJ/kg the gases leaving at an enthalpy of hm=616 Kj/kg here , for heating air the fuel requirement is "17 \frac{kg(air)}{kg(fuel)}" so,

Here it is given that hose required heat as 17.02 kW = "17.02 \frac{kJ}{s}"


Total heat requirement per day by hose will = "17.02\times3600\times24=1470528 kJ"


Net enthalpy exchange by air to heat the hose = "(h_f+h_a)-h_m=42731 kJ" /kg


so, net mass of air consumption per day "= \frac{1470528}{42731} =34.41 Kg" /day


so net mass of fuel consumption per day ="\frac{34.41}{17.02}=2.02 kg" of fuel per day